# Quantifying Lexical Semantic Shift via Unbalanced Optimal Transport

### v1

@@ Abstract
Ryo Kishino1   Hiroaki Yamagiwa1 Ryo Nagata2,4   Sho Yokoi3,4   Hidetoshi Shimodaira1,4 ${}^{1}\,$yoto University   ${}^{2}\,$onan University   ${}^{3}\,$ohoku University   ${}^{4}\,$ kishino.ryo.32s@st.kyoto-u.ac.jp, hiroaki.yamagiwa@sys.i.kyoto-u.ac.jp, nagata-acl2025@ml.hyogo-u.ac.jp, yokoi@tohoku.ac.jp, shimo@i.kyoto-u.ac.jp
@@ 1 Introduction
Lexical semantic change detection is the task of identifying words that change their meaning over time, as well as determining which specific senses have disappeared or emerged (Montanelli and Periti, 2023). Recently, methods leveraging word embeddings from a diachronic corpus pair have been studied. Previous approaches detect semantic change by measuring differences in contextualized embeddings of a target word across two corpora (Giulianelli et al., 2020; Montariol et al., 2021; Aida and Bollegala, 2023).
These methods focus on the entire set of usage instances. From a linguistic point of view, however, it is critical to focus on individual instances as well as the overall differences. For example, it is important to quantify the extent to which a disappearing or emerging sense of a word loses or acquires its popularity compared to the other senses in order to reveal how and why the word has undergone the semantic change. Unfortunately, there have been almost no studies on this topic.
@@ 2.1 Lexical semantic change detection
To address this, methods leveraging contextualized word embeddings for sets of usage instances of a target word, obtained from old and modern corpora using language models such as BERT (Devlin et al., 2019), have been proposed. These methods are broadly categorized into sense-based and form-based approaches (Giulianelli et al., 2020; Montanelli and Periti, 2023) as follows.
@@
This approach aims to estimate, either directly or indirectly, the frequency of word usage corresponding to specific senses of the target word in two given corpora. The estimated sense counts are then directly used to compute the degree of semantic change. Giulianelli et al. (2020); Rother et al. (2020); Kutuzov and Giulianelli (2020); Periti et al. (2022) applied clustering algorithms to two sets of contextualized embeddings of the target word to identify sense counts. Clustering methods such as $K$-means and Affinity Propagation (Frey and Dueck, 2007) are commonly used for this purpose. However, $K$-means faces the non-trivial challenge of determining the number of senses, while Affinity Propagation suffers from unstable clustering results (Montanelli and Periti, 2023).
@@
Unlike the sense-based approach, the form-based approach bypasses sense identification, which is likely to result in more stable semantic change detection. As mentioned in Eqs. (3) and (8) of Nagata et al. (2023), it is also possible to capture semantic change at the instance level by using the ratio of probability densities in the two corpora. While Nagata et al. (2023) discusses this approach, to the best of our knowledge, prior research has not quantitatively evaluated the detection of semantic change at the instance level.
@@ 3.1 Problem setting
We aim to identify how the meanings of a target word $w$ change across a diachronic corpus pair. A context containing the target word $w$ is referred to as a usage instance of $w$. For simplicity, it is referred to as an instance. The sense of $w$ in a usage instance is referred to as the sense of an instance. Let the set of $m$ usage instances of the target word $w$ belonging to the old corpus be $\{s^{w}_{i}\}_{i=1}^{m}$. Likewise, the set of $n$ usage instances of $w$ belonging to the modern corpus be $\{t^{w}_{j}\}_{j=1}^{n}$. Using a pre-trained language model, the contextualized embeddings $\bm{u}^{w}_{i}$ and $\bm{v}^{w}_{j}\in\mathbb{R}^{d}$ are calculated for each usage instance $s^{w}_{i}$ and $t^{w}_{j}$, respectively. Hereafter, the superscript $w$ is omitted for brevity. To detect how the meaning of $w$ has changed between the corpora, we measure the difference between the distributions $\{\bm{u}_{i}\}_{i=1}^{m}$ and $\{\bm{v}_{j}\}_{j=1}^{n}$.
@@ 3.2 Optimal transport
Let the weights associated with the contextualized word embeddings $\bm{u}_{i}$ and $\bm{v}_{j}$ be $a_{i},b_{j}\in\mathbb{R}_{+}$, respectively, satisfying $\sum_{i}a_{i}=\sum_{j}b_{j}=1$. Thus, these weighted embeddings can be regarded as probability distributions. Let $C_{ij}\in\mathbb{R}_{+}$ denote the transportation cost between the embeddings $\bm{u}_{i}$ and $\bm{v}_{j}$, and $T_{ij}\in\mathbb{R}_{+}$ the amount of transportation. The OT problem, which minimizes the total transportation cost, is formulated as follows:
@@ 3.3 Limitation of OT in semantic change detection
Figure 2: Illustration of the difference between OT and UOT. Circles and squares represent usage instances of the word record in the old and modern corpora, respectively, and the arrows indicate transportation. In the example above, the two bottom squares on the modern side represent instances in the newly emerged senses. (a) OT enforces a balanced alignment between the two sets of instances, which fails to model semantic change, as discussed in Section 3.3. (b) UOT allows for excess or deficit in alignment. The lack of transported weight from the old side to the two bottom squares on the modern side reflects the increased word usage of these new senses in the modern corpus, as detailed in Section 4.1.
@@ 4.1 Modeling semantic change with UOT
Here, $D_{1}:\mathbb{R}^{m}\times\mathbb{R}^{m}\to\mathbb{R}$ and $D_{2}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ are penalty functions, and $\lambda_{1},\lambda_{2}$ are hyperparameters that control the degree of penalization. The problem (2) is referred to as Unbalanced Optimal Transport (UOT). UOT allows for excess or deficit in the alignment by incurring a certain cost. See Appendix C for details of the UOT settings.
@@ 4.4 Log-density ratio
Figure 5: Re-rendering of the t-SNE visualization for the word record in Fig. 1. The color of each instance represents the Log-Density Ratio (LDR). For details and additional word examples, refer to Fig. 7 in Appendix A.
As a baseline for SUS in the form-based approach, we consider the log-density ratio (LDR). LDR is defined as the logarithm of the ratio of two probability density functions, which are parametrically estimated from the embeddings in the old and modern corpora. The estimation of density ratios has been extensively studied in statistics and machine learning (Sugiyama et al., 2012). In the context of semantic change detection, LDR is mentioned in Eq. (8) of Nagata et al. (2023), with further details provided in Appendix B. Estimating LDR in high-dimensional embedding spaces is challenging due to substantial fluctuations in the density ratio. SUS offers a non-parametric alternative, bypassing density ratio estimation via UOT.
A visual comparison of Fig. 1 and Fig. 5 suggests that both SUS and LDR roughly reflect sense clusters appropriately, although, as discussed in Appendix A, SUS performs slightly better than LDR. It is also worth noting that the range of fluctuations in LDR values is significantly large. A quantitative comparison between SUS and LDR in semantic change detection tasks will be presented later in the experimental sections.
@@ 5.1 Experimental configuration
In the experiments, we used the Diachronic Word Usage Graph (DWUG) dataset for English ver. 3 (Schlechtweg et al., 2021, 2024). It contains 46 target words. For each target word, about 100 usage instances are provided for an old time period (1810–1860) and also for a new time period (1960–2010). For details on DWUG, refer to Appendix D. In particular, an overview of the target word record is provided in Table 8 therein.
@@
We set uniform weights for each usage instance: $\bm{a}=\left(\frac{1}{m},\ldots,\frac{1}{m}\right)^{\top},\quad\bm{b}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)^{\top}$. We defined the transportation cost between instances using the cosine distance: $C_{ij}=1-\cos(\bm{u}_{i},\bm{v}_{j})$. For penalizing excess or deficit in the alignment, we used the L2 error: $D_{1}(\bm{T}\bm{1}_{n},\bm{a})=\|\bm{T}\bm{1}_{n}-\bm{a}\|_{2}^{2}$, $D_{2}(\bm{T}^{\top}\bm{1}_{m},\bm{b})=\|\bm{T}^{\top}\bm{1}_{m}-\bm{b}\|_{2}^{2}$. The MM algorithm provided by Python Optimal Transport (Flamary et al., 2021) was used for implementation. Following previous work on UOT (Chapel et al., 2021), we set $\lambda_{1}=\lambda_{2}=\lambda$. The value of $\lambda$ was kept consistent across all target words. For qualitative analysis in Section 4.3, $\lambda$ was fixed at $\lambda=100$. For quantitative evaluation using SUS in Sections 6 and 7, $\lambda$ was determined for each experiment by using a validation set.
@@ 5.2 Sense frequency distribution (SFD)
Here, we describe the gold standard [^3] The asterisk (*) in mathematical notation used in the experiments indicates the gold standard. used to evaluate the performance of the proposed methods and the baselines in the experiments. In the DWUG dataset, based on human annotations, each usage instance of a target word is computationally assigned a sense $k=1,\ldots,K^{*}$. Note that the senses themselves are not human-annotated, as detailed in Appendix D. In this study, we regard these sense labels in the DWUG dataset as the gold senses. Since the sense inventory is the same in the old and modern corpora, we can define the sense frequency distribution (SFD), representing the frequency of each sense $k$ in both corpora. Let $X_{k}^{*}$ and $Y_{k}^{*}$ denote the total number of instances associated with sense $k$ in the old and modern corpora, respectively. The SFDs are then expressed as $X^{*}=(X_{1}^{*},\ldots,X_{K^{*}}^{*})$ and $Y^{*}=(Y_{1}^{*},\ldots,Y_{K^{*}}^{*})$. Although the SFDs derived from DWUG may not always reflect the true SFDs, we assume them to be sufficiently reliable for our analysis.
@@ 6 Evaluation of Quantifying Instance-Level Semantic Change
Method Approach Spearman
$\tau_{\mathrm{SUS}}$ SUS-based 0.46
$\tau_{\mathrm{LDR}}$ form-based 0.40
$\tau_{\mathrm{WiDiD}}$ sense-based 0.31
Table 2: Performance of methods for predicting the instance-level change score.
Figure 6: The relationship between the gold change score $\tau^{*}$ derived from DWUG and the values of SUS (left) and LDR (right) for individual instances. The average values for each $\tau^{*}$ are marked with $\blacktriangle$. SUS exhibits a stronger Spearman correlation (0.85) with $\tau^{*}$ compared to LDR (0.74).
@@
Table 2 shows the Spearman rank correlation between the gold change scores $\tau^{*}$ and the change score produced by each method. SUS outperforms the other methods, providing more accurate predictions of changes in word usage frequency for the sense of each instance.
Fig. 6 shows the values of SUS and LDR for all usage instances of the target words on the horizontal axis and their corresponding gold change scores $\tau^{*}$ on the vertical axis. Since instances with the same sense share the same $\tau^{*}$, we compute the Spearman rank correlation between the gold change scores and the average SUS values for instances with the same $\tau^{*}$. SUS achieves a Spearman rank correlation of 0.85, indicating its effectiveness in capturing changes in usage frequency. In contrast, LDR shows a lower correlation of 0.74, demonstrating its inferior performance.

### v2

@@ Abstract
Ryo Kishino1   Hiroaki Yamagiwa1 Ryo Nagata2,5   Sho Yokoi3,4,5   Hidetoshi Shimodaira1,5 ${}^{1}\,$yoto University  ${}^{2}\,$onan University  ${}^{3}\,$  ${}^{4}\,$ohoku University  ${}^{5}\,$ kishino.ryo.32s@st.kyoto-u.ac.jp, h.yamagiwa@i.kyoto-u.ac.jp, nagata-acl2025@ml.hyogo-u.ac.jp, yokoi@ninjal.ac.jp, shimo@i.kyoto-u.ac.jp
@@ 1 Introduction
Lexical semantic change detection is the task of identifying words that change their meaning over time, as well as determining which specific senses have disappeared or emerged (Periti and Montanelli, 2024). Recently, methods leveraging word embeddings from a diachronic corpus pair have been studied. Previous approaches detect semantic change by measuring differences in contextualized embeddings of a target word across two corpora (Giulianelli et al., 2020; Montariol et al., 2021; Aida and Bollegala, 2023).
These methods focus on the entire set of usage instances. From a linguistic point of view, however, it is critical to focus on individual instances as well as on the overall differences. For example, it is important to quantify the extent to which a disappearing or emerging sense of a word loses or acquires its popularity compared to the other senses in order to reveal how and why the word has undergone the semantic change. Unfortunately, there have been almost no studies on this topic.
@@ 2.1 Lexical semantic change detection
To address this, methods leveraging contextualized word embeddings for sets of usage instances of a target word, obtained from old and modern corpora using language models such as BERT (Devlin et al., 2019), have been proposed. These methods are broadly categorized into sense-based and form-based approaches (Giulianelli et al., 2020; Periti and Montanelli, 2024) as follows.
@@
This approach aims to estimate, either directly or indirectly, the frequency of word usage corresponding to specific senses of the target word in two given corpora. The estimated sense counts are then directly used to compute the degree of semantic change. Giulianelli et al. (2020); Rother et al. (2020); Kutuzov and Giulianelli (2020); Periti et al. (2022) applied clustering algorithms to two sets of contextualized embeddings of the target word to identify sense counts. Clustering methods such as $K$-means and Affinity Propagation (Frey and Dueck, 2007) are commonly used for this purpose. However, $K$-means faces the non-trivial challenge of determining the number of senses, while Affinity Propagation suffers from unstable clustering results (Periti and Montanelli, 2024).
@@
Unlike the sense-based approach, the form-based approach bypasses sense identification, which is likely to result in more stable semantic change detection. As mentioned in Eqs. (3) and (8) of Nagata et al. (2023), it is also possible to capture semantic change at the instance level by using the ratio of probability densities in the two corpora. While Nagata et al. (2023) discuss this approach, to the best of our knowledge, prior research has not quantitatively evaluated the detection of semantic change at the instance level.
@@ 3.1 Problem setting
We aim to identify how the meanings of a target word $w$ change across a diachronic corpus pair. A context containing the target word $w$ is referred to as a usage instance of $w$. For simplicity, it is referred to as an instance. The sense of $w$ in a usage instance is referred to as the sense of an instance. Let the set of $m$ usage instances of the target word $w$ belonging to the old corpus be $\{s^{w}_{i}\}_{i=1}^{m}$. Likewise, let the set of $n$ usage instances of $w$ belonging to the modern corpus be $\{t^{w}_{j}\}_{j=1}^{n}$. Using a pre-trained language model, the contextualized embeddings $\bm{u}^{w}_{i}$ and $\bm{v}^{w}_{j}\in\mathbb{R}^{d}$ are calculated for each usage instance $s^{w}_{i}$ and $t^{w}_{j}$, respectively. Hereafter, the superscript $w$ is omitted for brevity. To detect how the meaning of $w$ has changed between the corpora, we measure the difference between the distributions $\{\bm{u}_{i}\}_{i=1}^{m}$ and $\{\bm{v}_{j}\}_{j=1}^{n}$.
@@ 3.2 Optimal transport
Let $a_{i}\in\mathbb{R}_{+}$ and $b_{j}\in\mathbb{R}_{+}$ denote the weights associated with the contextualized word embeddings $u_{i}$ and $v_{j}$, respectively. satisfying $\sum_{i}a_{i}=\sum_{j}b_{j}=1$. Thus, these weighted embeddings can be regarded as probability distributions. Let $C_{ij}\in\mathbb{R}_{+}$ denote the transportation cost between the embeddings $\bm{u}_{i}$ and $\bm{v}_{j}$, and $T_{ij}\in\mathbb{R}_{+}$ the amount of transportation. The OT problem, which minimizes the total transportation cost, is formulated as follows:
@@ 3.3 Limitation of OT in semantic change detection
Figure 2: Illustration of the difference between OT and UOT. Circles and squares represent usage instances of the word record in the old and modern corpora, respectively, and the arrows indicate transportation. In the example above, the two bottom squares on the modern side represent instances in the newly emerged senses. (a) OT enforces a balanced alignment between the two sets of instances, which fails to model semantic change, as discussed in Section 3.3. (b) UOT allows for excess or deficit in alignment. The lack of transported weight from the old side to the two bottom squares on the modern side reflects the increased word usage of these new senses in the modern corpus, as described in Section 4.1. See Fig. 3 for a visualization of the transport matrix based on actual numerical experiments.
@@ 4.1 Modeling semantic change with UOT
Here, $D_{1}:\mathbb{R}^{m}\times\mathbb{R}^{m}\to\mathbb{R}$ and $D_{2}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ are penalty functions, and $\lambda_{1},\lambda_{2}$ are hyperparameters that control the degree of penalization. The problem (2) is referred to as Unbalanced Optimal Transport (UOT). UOT allows for excess or deficit in the alignment by incurring a certain cost. In our experiments, we assigned uniform weights to each embedding and employed the L2 error as the penalty function. For details, see Section 5.
@@ 4.4 Log-density ratio
(a) LDR
(b) DWUG
Figure 5: Re-rendering of the t-SNE visualization for the word record in Fig. 1. The color of each instance represents (a) the Log-Density Ratio (LDR) and (b) the gold sense in the DWUG dataset. For details and additional word examples with LDR coloring, refer to Fig. 7 in Appendix A. For details on the gold sense, see Section 5.2.
Definition and computation.
As a baseline for SUS in the form-based approach, we consider the log-density ratio (LDR). LDR is defined as the logarithm of the ratio of two probability density functions, which are parametrically estimated from the embeddings in the old and modern corpora. The estimation of density ratios has been extensively studied in statistics and machine learning (Sugiyama et al., 2012).
In the context of semantic change detection, LDR is mentioned in Eq. (8) of Nagata et al. (2023). Note that they use the term log-likelihood ratio to refer to the LDR at a single instance, although it typically refers to comparisons between densities over the entire dataset.
In our setup, we compute LDR as follows. First, following Nagata et al. (2023), we assume that the normalized old and modern embeddings follow a distinct von Mises-Fisher (vMF) distribution [^3] The probability density function of the vMF distribution with mean direction parameter $\bm{\mu}$ and concentration parameter $\kappa$ is given by $p(\bm{x}\mid\bm{\mu},\kappa)\propto\exp(\kappa\bm{\mu}^{\top}\bm{x})$, where $\|\bm{x}\|=1$. , $p_{S}$ and $p_{T}$. Then, we perform maximum likelihood estimation of the parameters using an approximate closed-form solution, as described in Appendix B. Using the estimated parameters, we compute
$\mathrm{LDR}(s_{i})$ $=\log\frac{p_{T}(\tilde{\bm{u}}_{i})}{p_{S}(\tilde{\bm{u}}_{i})},$
$\mathrm{LDR}(t_{j})$ $=\log\frac{p_{T}(\tilde{\bm{v}}_{j})}{p_{S}(\tilde{\bm{v}}_{j})},$
where $\tilde{\bm{u}}_{i}=\bm{u}_{i}/\|\bm{u}_{i}\|$ and $\tilde{\bm{v}}_{j}=\bm{v}_{j}/\|\bm{v}_{j}\|$.
Comparison between SUS and LDR.
A visual comparison of Fig. 1 and Fig. 5 suggests that both SUS and LDR roughly reflect sense clusters appropriately. However, referring to the gold senses [^4] In the dataset used in the experiments, word senses are defined automatically based on human annotations. See Section 5.2 for further details. provided by the DWUG dataset in Fig. 5, it is evident that SUS reflects the clusters more accurately than LDR. In particular, in the rightmost part of the cluster corresponding to the sense [information], there is a red region where circles and boxes are mixed. LDR tends to separate this region as a distinct cluster, although this separation does not match the gold senses. SUS, on the other hand, better preserves the overall cluster structure. For preprocessing on color configuration, see Appendix A.
LDR values often fluctuate substantially, making their estimation in high-dimensional spaces unstable. SUS offers a non-parametric alternative that bypasses this instability via UOT.
A detailed comparison between SUS and LDR in semantic change detection tasks will be presented later in Sections 6 and 7.
@@ 5.1 Experimental configuration
In the experiments, we used the Diachronic Word Usage Graph (DWUG) dataset for English version 3 (Schlechtweg et al., 2021, 2024). It contains 46 target words. For each target word, about 100 usage instances are provided for an old time period (1810–1860) and also for a new time period (1960–2010). For details on DWUG, refer to Appendix D. In particular, an overview of the target word record is provided in Table 8 therein.
In addition, to verify that the proposed method does not overfit to a single dataset, we also conducted experiments on DWUG ES (Zamora-Reina et al., 2022), a dataset designed to detect semantic change in Spanish. This dataset was constructed following the same procedure used to construct English DWUG dataset. The results on this dataset are presented in Appendix H.
@@
We set uniform weights for each usage instance: $\bm{a}=\left(\frac{1}{m},\ldots,\frac{1}{m}\right)^{\top},\quad\bm{b}=\left(\frac{1}{n},\ldots,\frac{1}{n}\right)^{\top}$. We defined the transportation cost between instances using the cosine distance: $C_{ij}=1-\cos(\bm{u}_{i},\bm{v}_{j})$. For penalizing excess or deficit in the alignment, we used the L2 error: $D_{1}(\bm{T}\bm{1}_{n},\bm{a})=\|\bm{T}\bm{1}_{n}-\bm{a}\|_{2}^{2}$, $D_{2}(\bm{T}^{\top}\bm{1}_{m},\bm{b})=\|\bm{T}^{\top}\bm{1}_{m}-\bm{b}\|_{2}^{2}$. The majorization-minimization (MM) algorithm provided by Python Optimal Transport (Flamary et al., 2021) was used for implementation. Following previous work on UOT (Chapel et al., 2021), we set $\lambda_{1}=\lambda_{2}=\lambda$. The value of $\lambda$ was kept consistent across all target words. For qualitative analysis in Section 4.3, $\lambda$ was fixed at $\lambda=100$. For quantitative evaluation using SUS in Sections 6 and 7, $\lambda$ was determined for each experiment by using a validation set.
Regarding the weights $\bm{a}$ and $\bm{b}$, one alternative is to use the norms of word embeddings, as proposed by Yokoi et al. (2020). However, this choice had little impact on performance in the task described in Section 7. This is likely because all embeddings represent the same target word, and their norms do not vary substantially.
As a penalty function, the Kullback-Leibler divergence has also been used in UOT (Chapel et al., 2021; Arase et al., 2023), but we consider that using the L2 error is more natural for focusing on the excess and deficit of alignment.
@@ 5.2 Sense frequency distribution (SFD)
Here, we describe the gold standard [^5] The asterisk (*) in mathematical notation used in the experiments indicates the gold standard. used to evaluate the performance of the proposed methods and the baselines in the experiments. In the DWUG dataset, based on human annotations, each usage instance of a target word is computationally assigned a sense $k=1,\ldots,K^{*}$. Note that the senses themselves are not human-annotated, as described in Appendix D. In this study, we regard these sense labels in the DWUG dataset as the gold senses. Since the sense inventory is the same in the old and modern corpora, we can define the sense frequency distribution (SFD), representing the frequency of each sense $k$ in both corpora. Let $X_{k}^{*}$ and $Y_{k}^{*}$ denote the total number of instances associated with sense $k$ in the old and modern corpora, respectively. The SFDs are then expressed as $X^{*}=(X_{1}^{*},\ldots,X_{K^{*}}^{*})$ and $Y^{*}=(Y_{1}^{*},\ldots,Y_{K^{*}}^{*})$. Although the SFDs derived from DWUG may not always reflect the true SFDs, we assume them to be sufficiently reliable for our analysis.
@@ 6 Evaluation of Quantifying Instance-Level Semantic Change
Method Approach Instance-level Sense-level
$\tau_{\mathrm{SUS}}$ SUS-based 0.46 0.83
$\tau_{\mathrm{LDR}}$ form-based 0.40 0.70
$\tau_{\mathrm{WiDiD}}$ sense-based 0.31 0.84
Table 2: Performance of methods for predicting the instance-level and sense-level change score.
Figure 6: The relationship between the gold change score $\tau^{*}$ derived from DWUG and the values of SUS (left) and LDR (right) for all instances. The average values for each $\tau^{*}$ are marked with $\blacktriangle$. The Spearman rank correlation computed over $\bullet$ corresponds to instance-level performance and $\blacktriangle$ to sense-level performance.
@@
Table 2 shows the Spearman rank correlation between the gold change scores $\tau^{*}$ and the change score produced by each method. To clarify the situation, see Fig. 6, which shows the values of SUS and LDR for all usage instances of the target words on the horizontal axis and their corresponding gold change scores $\tau^{*}$ on the vertical axis. SUS outperforms the other methods, providing more accurate predictions of changes in word usage frequency for the sense of each instance. Since instances with the same sense share the same $\tau^{*}$, we compute the Spearman rank correlation between the gold change scores and the average SUS values for instances with the same $\tau^{*}$. We call this sense-level semantic change. SUS achieves a Spearman rank correlation of 0.83, indicating its effectiveness in capturing changes in usage frequency. In contrast, LDR shows a lower correlation of 0.70, demonstrating its inferior performance. While WiDID performs well for sense-level semantic change detection, it is worth noting that the gold scores are based on sense-level semantic change, which inherently favors WiDID as a metric.